Emmanuel is given a multiple-choice exam with ten questions and each question with
fi ve possible answers. He decided to guess randomly for each question. What is the
probability that he will get at least six questions correct?
1
Expert's answer
2020-06-01T19:51:49-0400

Let "X=" the number of correctly solved questions: "X\sim Bin (n,p)."


"P(X=x)=\binom{n}{x}p^x(1-p)^{n-x}"

Given "n=10,p=0.2."


"P(X\geq 6)=P(X=6)+P(X=7)+P(X=8)+"

"P(X=9)+P(X=10)=\binom{10}{6}(0.2)^6(1-0.2)^{10-6}+"

"+\binom{10}{7}(0.2)^7(1-0.2)^{10-7}+\binom{10}{8}(0.2)^8(1-0.2)^{10-8}+"

"+\binom{10}{9}(0.2)^9(1-0.2)^{10-9}+\binom{10}{10}(0.2)^{10}(1-0.2)^{10-10}="

"=210(0.2)^6(0.8)^4+120(0.2)^7(0.8)^3+45(0.2)^8(0.8)^2+"

"+10(0.2)^9(0.8)+(0.2)^{10}="

"=0.005505024+0.000786432+0.000073728+"

"+0.000004096+0.0000001024\approx0.0064"